OpenStudy (anonymous):

What are the coordinates of the vertex of the parabola given by the equation: y = x2 - 2x - 2 ? A. (0, -2) B. (1, -3) C. (4, 6) D. (-2, 6)

OpenStudy (ranga):

The general equation of a parabola is: y = ax^2 + bx + c ---- (1) The x-coordinate of the vertex of the parabola is: -b/(2a) Once you find the x-coordinate you can find the y-coordinate by putting the x value in (1)

OpenStudy (anonymous):

at vertex of the parabola slope of graph is Zero. so lets first calculate slope and put that equal to zero. \[\frac{ dy }{ dx }= 2x-2 = 0\] \[\implies x = 1\] when \[x=1\], \[y =-3\] hence vertex is (1,-3)